Express 2log 3 + 3log 2 - log 6 as a single logarithm.
Answer by Alan3354(69443)

Express 2log 3 + 3log 2 - log 6 as a single logarithm.
-----
2log(3) = log(9)
3log(2) = log(8)
----
--> log(8*9/6)
= log(12)